# Boson-fermion pairing and condensation in two-dimensional Bose-Fermi mixtures

### v1

@@ IV.6 Identification of different physical regions in the coupling-concentration plane
Specifically, in Sec. III two different regimes were identified when varying the bosonic concentration for a fixed strong BF attraction. Analogously, two different regimes were identified in Sec. IV.2 when varying the coupling strength. Gathering all this information together, a schematic picture can be constructed as in Fig. 16, whereby a number of different physical regimes are detected:
  • Weakly interacting BF mixture: a majority of weakly perturbed Fermi atoms coexists with a mildly depleted bosonic condensate; bosons out of the condensate marginally engage in molecule formation;
@@ V Conclusions and outlook
A further difference with respect to the 3D case is the hybrid nature of the composite fermions that form for sufficiently strong BF attraction, with associated rather unusual momentum distributions, except for the case $x=1$ in which the mixture “fermionizes” completely in a gas of non-interacting point-like fermionic molecules, similarly to what is obtained in 3D for all $x\leq 1$. The origin of this difference for $x<1$ is in the residual condensate density $n_{0}$ which, albeit small, proves sufficient to generate a significant hybridization through the hybridization energy-scale $\Delta_{0}\propto(n_{0}{\varepsilon_{0}})^{1/2}$.
A discussion on possible refinements of the present approach is finally in order. The absence of a quantum phase transition for the condensate fraction found in the present work, recalls the absence of the polaron-to-molecule transition that is found in the single impurity limit of a Fermi polaron, when the polaron is described by a non-self-consistent $T$-matrix self-energy [109]. In this limit, it has been pointed out that a correct description of the atom-dimer interaction is crucial in bringing about this transition [109, 110].
In Sec. III.1 and Sec. III.2, we mentioned that the non-self-consistent $T$-matrix approximation takes into account the atom-dimer interaction at the level of the Born approximation, resulting in coupling-independent mean-field shifts $\Sigma_{\mathrm{CF}}$ and $\Sigma^{0}_{\mathrm{F}}$. While in 3D the Born approximation is qualitatively correct, in 2D these Hartree-like contributions definitely overestimate the effective repulsion between molecules and unpaired fermionic atoms, making the molecular state energetically less convenient, and favoring a state in which the molecule hybridizes with an unpaired state made of a fermionic atom and a boson belonging to the condensate. In order to go beyond the Born approximation, one should, in principle, include in a many-body theory the same kind of diagrams (involving two fermions repeatedly exchanging a boson) describing the atom-dimer scattering in the three-body problem [111, 112, 97] (see also [113]), a task that is far from trivial.
Insight from the polaron limit may suggest that the absence of a critical coupling strength in 2D is simply an artifact of the approximation for the self-energy here adopted. We do not exclude such a possibility. We however stress that, in experiments, the exponential vanishing of the condensate at large coupling strength, which we have found in the present work, would hardly be distinguishable from an exact vanishing of the condensate (leaving aside the fact that at finite temperature, as experimentally relevant, one should consider a quasi-condensate in 2D, introducing a further source of smearing of the transition).
@@
Finally, it is worth pointing out that, since the polaron-to-molecule quantum phase transition is proven to occur specifically in the impurity limit, one may expect that in the high concentration regime, where the polaron picture breaks down, this transition could be suppressed or significantly modified by quantum fluctuations. It is known in 3D [114, 33] that, with a finite density of impurities, thermal and quantum fluctuations smoothens the discontinuity arising in the single impurity limit thus shifting the critical coupling to higher values. An analogous and stronger role may be played by quantum fluctuations in 2D. To settle this point, a detailed calculation of the three-body diagrammatic contributions for finite densities beyond the polaronic regime, should be addressed in future work.
Acknowledgements.
We thank Dr. Andrea Guidini for contributing in an early stage of this work. We acknowledge the use of computational resources from the parallel computing cluster of the Open Physics Hub at the Physics and Astronomy Department in Bologna. Financial support from the Italian Ministry of University and Research under project PRIN2022, Contract No. 2022523NA7 and from the European Union - NextGenerationEU through the Italian Ministry of University and Research under PNRR - M4C2 - I1.4 Contract No. CN00000013 is also acknowledged.
Appendix A Normal phase $T$-matrix $\Gamma$(${\mathbf{P}}$,$\Omega$)
Evaluating the integral over $\omega$ in Eq. (3) on the complex plane by closing the contour on the left-hand side and rearranging the terms of the resulting geometric series, one obtains
@@ Appendix A Normal phase $T$-matrix $\Gamma$(${\mathbf{P}}$,$\Omega$)
whereby the $\Theta$ functions appearing in the integrand descend from the Fermi and Bose distributions in the zero temperature limit. In particular, for the coupling strengths and concentrations considered in this work, ${\mu_{\mathrm{B}}}$ is always negative, hence $\Theta(-\xi_{\mathbf{k}}^{\mathrm{B}})=0$.
@@
$\Gamma\left(\textbf{P},\Omega\right)^{-1}=T_{2}({\mathbf{P}},\Omega)^{-1}-I_{\mathrm{F}}({\mathbf{P}},\Omega),$ (65)
where $T_{2}$ is the off-shell two-body $T$-matrix in vacuum [93], which for $\Omega\neq 0$ is given by
@@
$=$ $-\frac{m_{r}}{2\pi}\ln\left(\frac{\frac{P^{2}}{2M}-\mu_{\mathrm{F}}-\mu_{\mathrm{B}}-i\Omega}{{\varepsilon_{0}}}\right).$
@@
$I_{\mathrm{F}}\left(\textbf{P},\Omega\right)=\int\frac{d{\mathbf{k}}}{\left(2\pi\right)^{2}}\frac{\Theta\left(-\xi_{\textbf{P}-{\mathbf{k}}}^{\mathrm{F}}\right)}{\xi_{\textbf{P}-{\mathbf{k}}}^{\mathrm{F}}+\xi^{\mathrm{B}}_{\mathbf{k}}-i\Omega},$ (67)
@@
For $\mu_{\mathrm{F}}>0$ Eq. (67) can be expressed in closed form as follows (with $\Omega\neq 0$)
$I_{\mathrm{F}}({\mathbf{P}},\Omega)=\frac{m_{r}}{2\pi}\begin{dcases}\ln\frac{\mathcal{A}}{\alpha+\beta}\quad&\text{for }{\mu_{\mathrm{F}}}+{\mu_{\mathrm{B}}}\leq\frac{P^{2}}{2m_{\mathrm{B}}};\\ \ln{\mathcal{A}\left(\alpha-\beta\right)\over\mathcal{B}^{2}}+i\pi\text{sgn}(\Omega)\quad&\text{for }{\mu_{\mathrm{F}}}+{\mu_{\mathrm{B}}}>\frac{P^{2}}{2m_{\mathrm{B}}}\text{ and }k_{+}\leq k_{\mu_{\mathrm{F}}};\\ -\ln\frac{\mathcal{A}}{\alpha-\beta}+i\pi\text{sgn}\left(\Omega\right)\quad&\text{for }{\mu_{\mathrm{F}}}+{\mu_{\mathrm{B}}}>\frac{P^{2}}{2m_{\mathrm{B}}}\text{ and }k_{+}>k_{\mu_{\mathrm{F}}},\end{dcases}$ (68)
@@
$\mathcal{A}$ $=\beta+\frac{k^{2}_{\mu_{\mathrm{F}}}}{2m_{r}}+\sqrt{\alpha^{2}+\beta{k^{2}_{\mu_{\mathrm{F}}}\over m_{r}}+{k^{4}_{\mu_{\mathrm{F}}}\over 4m_{r}^{2}}},$ (69)
$\mathcal{B}^{2}$ $=\left(i\Omega+\frac{P^{2}}{m_{\mathrm{B}}^{2}}m_{r}\right)^{2}+\frac{P^{2}}{m_{\mathrm{B}}^{2}}k_{+}^{2}+\Omega^{2},$ (70)
@@
${k_{+}^{2}\over 2m_{r}}$ $={\mu_{\mathrm{F}}}+{\mu_{\mathrm{B}}}-\frac{P^{2}}{2m_{\mathrm{B}}},\quad\quad{k_{\mu_{\mathrm{F}}}^{2}\over 2m_{\mathrm{F}}}=\mu_{\mathrm{F}},$
$\alpha$ $=-{k_{+}^{2}\over 2m_{r}}-i\Omega,\quad\quad\quad\quad\beta=\alpha-\frac{P^{2}}{m_{\mathrm{B}}^{2}}m_{r}.$
@@
We briefly illustrate the evaluation of the 3-body $T$-matrix within Born approximation in the 2D vacuum. We start by recalling the Skorniakov-Ter-Martirosian equation for the 3-body $T$-matrix $T_{3}(k,p;K)$ describing the scattering between a BF dimer and a fermionic atom of initial three-momenta $K-k$ and $k$ and final three-momenta $K-p$ and $p$, respectively [115, 111]:
$T_{3}(k,p;K)=-G_{\mathrm{B}}^{0}(K-k-p)$
$-\int\!\!\!\frac{d\mathbf{q}}{(2\pi)^{2}}\frac{dq_{0}}{(2\pi)}G_{\mathrm{B}}^{0}(K-k-q)G_{\mathrm{F}}^{0}(q)T_{2}(K-q)T_{3}(q,p;K),$ (71)
@@
$G_{s}^{0}(q)=\frac{1}{q_{0}-\mathbf{q}^{2}/2m_{s}+i0^{+}}\quad\quad s=\mathrm{B,F}.$ (72)
Within Born approximation, $T_{3}$ is replaced with the first term on the right-hand side of Eq. 71. Moreover, for low energy scattering one can set $k_{1}\simeq k_{2}\simeq 0$ and $K\equiv\{\textbf{0},-{\varepsilon_{0}}\}$, (taking into account the binding energy ${\varepsilon_{0}}$ of the BF dimer). Recalling that $T_{2}$ is a dimer propagator except for the form factor $2\pi{\varepsilon_{0}}/m_{r}$ (see Eq. 19) due to its composite nature, the same normalization factor is applied to $T_{3}$, which thus becomes independent of the two-body scattering length $a_{\mathrm{BF}}$
${2\pi{\varepsilon_{0}}\over m_{r}}T_{3}(0,0;\{\textbf{0},-{\varepsilon_{0}}\})\simeq-{2\pi{\varepsilon_{0}}\over m_{r}}G_{\mathrm{B}}^{0}(\textbf{0},-{\varepsilon_{0}})={2\pi\over m_{r}}.$ (73)

### v2

@@ 4.6 Identification of different physical regions in the coupling-concentration plane
Specifically, in Sec. 3.2 and  3.4 two different regimes were identified when varying the bosonic concentration for a fixed strong BF attraction. Analogously, two different regimes were identified in Sec. 4.3 when varying the coupling strength, while the analysis of the condensate fraction of Sec. 4.4 added further insight into the evolution with the coupling strength.
Gathering all this information together, a schematic picture can be constructed as in Fig. 16, whereby a number of different physical regimes are detected:
  • Weakly interacting BF mixture: a majority of weakly perturbed Fermi atoms coexist with a mildly depleted bosonic condensate; bosons out of the condensate marginally engage in molecule formation;
@@ 5 Conclusions and outlook
A further difference with respect to the 3D case is the hybrid nature of the composite fermions that form for sufficiently strong BF attraction. Rather unusual features are found within the momentum distribution function, except for the case $x=1$ in which the mixture “fermionizes” completely in a gas of non-interacting point-like fermionic molecules, similarly to what is obtained in 3D for all $x\leq 1$. The origin of this difference for $x<1$ is in the residual condensate density $n_{0}$ which, albeit small, proves sufficient to generate a significant hybridization through the hybridization energy-scale $\Delta_{0}\propto(n_{0}{\varepsilon_{0}})^{1/2}$.
A discussion on possible refinements of the present approach is finally in order. The absence of a quantum phase transition for the condensate fraction found in the present work recalls the absence of the polaron-to-molecule transition that is found in the single impurity limit of a Fermi polaron when the polaron is described by a non-self-consistent $T$-matrix self-energy [97]. In this limit, it has been pointed out that a correct description of the atom-dimer interaction is crucial in bringing about this transition [97, 98].
In Sec. 3.1 and Sec. 3.2, we mentioned that the non-self-consistent $T$-matrix approximation takes into account the atom-dimer interaction at the level of the Born approximation, resulting in coupling-independent mean-field shifts $\Sigma^{0}_{\mathrm{CF}}$ and $\Sigma^{0}_{\mathrm{F}}$. While in 3D the Born approximation is qualitatively correct, in 2D these Hartree-like contributions definitely overestimate the effective repulsion between molecules and unpaired fermionic atoms, making the molecular state energetically less convenient, and favoring a state in which the molecule hybridizes with an unpaired state made of a fermionic atom and a boson belonging to the condensate. In order to go beyond the Born approximation, one should, in principle, include in a many-body theory the same kind of diagrams (involving two fermions repeatedly exchanging a boson) describing the atom-dimer scattering in the three-body problem [124, 125, 108] (see also [126]), a task that is far from trivial.
Insight from the polaron limit may suggest that the absence of a critical coupling strength in 2D is simply an artifact of the approximation for the self-energy here adopted. We do not exclude such a possibility, in particular at small values of $x$ which are closer to the single impurity limit. We however stress that, in experiments, the exponential vanishing of the condensate at large coupling strength, which we have found in the present work, would hardly be distinguishable from an exact vanishing of the condensate (leaving aside the fact that at finite temperature, as experimentally relevant, one should consider a quasi-condensate in 2D, introducing a further source of smearing of the transition).
@@
In addition, we point out that, since the polaron-to-molecule quantum phase transition is proven to occur specifically in the impurity limit, one cannot exclude that in the high concentration regime, where the polaron picture breaks down, this transition could be suppressed or significantly modified by quantum fluctuations. It is known that in 3D, with a finite density of impurities, thermal and quantum fluctuations smoothen the discontinuity arising in the single impurity limit thus shifting the critical coupling to higher values [127, 33]. An analogous and stronger role may be played by quantum fluctuations in 2D. To settle this point, a detailed calculation of the three-body diagrammatic contributions for finite densities beyond the polaronic regime should be addressed in future work.
It is also worth discussing the relevance of the present investigation to BF mixtures realized with transition metal dichalcogenides. The first issue is whether the effective interaction between excitons and excess charge carriers (the bosons and fermions of the mixture, respectively) can be modeled by a contact potential of the type utilized in the present work. In TMD systems, excitons and charge carriers can bind into a molecular trion state, which controls the effective BF interaction. The binding energy of the trion is typically at least one order of magnitude smaller than the exciton binding energy [128], and of the order of the Fermi energy of the excess charge carrier. Under these conditions, it has been shown that the effective BF interaction can indeed be reasonably modeled by a contact interaction [129]. A difference with respect to the model discussed in the present paper may instead be represented by the presence of Coulomb interactions between fermions, in particular intra-species interactions between excess charge carriers. It can, however, be argued that [130], provided the doping is not too low [90, 56] (in order to be away from the regime of Wigner crystallization and reduced screening), their effect will be mostly to renormalize the single-particle dispersions of excess charges, an effect that can be safely described by the parameters of Fermi liquid theory (notably, by the effective mass of the charges). Concerning, finally, exciton-exciton interaction, it is expected to be small due to the tightly-bound nature of the excitons in TMDs that makes their dipole moment small, with a sign of the interaction that depends on the orientation of the dipoles. We have already mentioned that the mechanical stability of BF mixtures requires a repulsive interaction. In this respect, bilayer TMDs with interlayer excitons would be preferable, since in this case the dipole moments will be essentially parallel to each other (we notice that the distance between layers[42] is comparable with the exciton radius [90]), thus ensuring the interaction to be repulsive. It is however not clear if such a repulsion would be sufficient to guarantee stability.
As a final remark, we wish to comment on possible extensions of the present work to finite temperature. In this case, the $T=0$ condensed phase with long-range order will be replaced by a BKT superfluid phase with quasi-long-range order for the bosonic component. Similarly to $T=0$, we expect BF pairing to compete with the bosonic superfluid phase. On a technical side, the main difficulty will be how to effectively include the BKT physics of the bosonic component in our diagrammatic scheme.
All numerical data necessary to reproduce Figs. 3-15 are available online [131].
Acknowledgements
We thank Dr. Andrea Guidini for contributing in an early stage of this work. We acknowledge the use of the parallel computing cluster of the Open Physics Hub at the Department of Physics and Astronomy of the University of Bologna.
Funding information
L.P.  and P.P.  acknowledge financial support from the Italian Ministry of University and Research (MUR) under project PRIN2022, Contract No. 2022523NA7. P.P. also acknowledges financial support from MUR project PE0000023-NQSTI (Italy) and from European Union - NextGenerationEU through the Italian Ministry of University and Research (MUR) under PNRR - M4C2 - I1.4 Contract No. CN00000013.
Appendix A Particle-particle ladder $\Gamma$(${\mathbf{P}}$, $\Omega$)
We start from Eq. (5) for $\Gamma({\bf P},\Omega)$. For $\mu_{\rm B}\leq 0$, it corresponds to
@@ Appendix A Particle-particle ladder $\Gamma$(${\mathbf{P}}$, $\Omega$)
$=$ $T_{2}({\mathbf{P}},\Omega)^{-1}-I_{\mathrm{F}}({\mathbf{P}},\Omega),$ (A.2)
where $T_{2}$ is the off-shell two-body $T$-matrix in vacuum [104] which, for $\Omega\neq 0$, is given by
@@
$=$ $-\frac{m_{r}}{2\pi}\ln\left(\frac{\frac{P^{2}}{2M}-\mu_{\mathrm{F}}-\mu_{\mathrm{B}}-i\Omega}{{\varepsilon_{0}}}\right).$ (A.4)
@@
$I_{\mathrm{F}}\left(\textbf{P},\Omega\right)=\int\!\!\!\frac{d{\mathbf{k}}}{\left(2\pi\right)^{2}}\frac{\Theta\left(-\xi_{\textbf{P}-{\mathbf{k}}}^{\mathrm{F}}\right)}{\xi_{\textbf{P}-{\mathbf{k}}}^{\mathrm{F}}+\xi^{\mathrm{B}}_{\mathbf{k}}-i\Omega},$ (A.5)
@@
For $\mu_{\mathrm{F}}>0$ and $\Omega\neq 0$, Eq. (A.5) can be expressed in closed form as
$I_{\mathrm{F}}({\mathbf{P}},\Omega)=\frac{m_{r}}{2\pi}\left[\ln\left(\mathcal{A}\right)-\ln\left(\frac{P^{2}}{2M}-\mu_{\mathrm{F}}-\mu_{\mathrm{B}}-i\Omega\right)-i\pi\textrm{sgn}\left(\Omega\right)\right]$ (A.6)
@@
$\mathcal{A}=\frac{z}{2}+\frac{P^{2}}{2M\gamma_{m}}+\textrm{sgn}\left(\textrm{Re}\left[z\right]\right)\sqrt{\left(\frac{z}{2}\right)^{2}-\frac{P^{2}}{2{m_{\mathrm{B}}}\gamma_{m}}{\mu_{\mathrm{F}}}},$ (A.7)
@@
$z={\mu_{\mathrm{B}}}-\frac{{\mu_{\mathrm{F}}}}{\gamma_{m}}-\frac{P^{2}}{2{m_{\mathrm{B}}}}+i\Omega,\quad\quad\quad\gamma_{m}=\frac{{m_{\mathrm{B}}}}{{m_{\mathrm{F}}}}.$ (A.8)
Note that, when inserted in Eq. (A.2), the second term in Eq. (A.6) perfectly cancels the logarithm of the numerator of Eq. (A.4), yielding
$\Gamma\left({\mathbf{P}},\Omega\right)=\frac{2\pi/m_{r}}{\ln\left(\varepsilon_{0}/\mathcal{A}\right)+i\pi\textrm{sgn}\left(\textrm{Im}\left[\mathcal{A}\right]\right)}.$ (A.9)
For equal masses and real frequencies, an equation of the same form as Eq.  (A.9) was first obtained in [121].
Finally, we derive strong-coupling expressions for $T_{2}({\mathbf{P}},\Omega)$ and $I_{\rm F}({\mathbf{P}},\Omega)$. For $T_{2}({\mathbf{P}},\Omega)$, we first write $\mu_{\rm F}+\mu_{\rm B}=\mu_{\rm CF}-{\varepsilon_{0}}$ in Eq. (A.2) and obtain
$T_{2}({\mathbf{P}},\Omega)^{-1}$ $=$ $-\frac{m_{r}}{2\pi}\ln\left(1+\frac{\frac{P^{2}}{2M}-\mu_{\mathrm{CF}}-i\Omega}{{\varepsilon_{0}}}\right)$ (A.10)
$\approx$ $\frac{m_{r}}{2\pi{\varepsilon_{0}}}\left(i\Omega-\frac{P^{2}}{2M}+\mu_{\mathrm{CF}}\right),$ (A.11)
where we have assumed $\Omega,P^{2}/2M,\mu_{\rm CF}\ll{\varepsilon_{0}}$ in the strong-coupling limit ${\varepsilon_{0}}\to\infty$.
For $I_{\rm F}({\mathbf{P}},\Omega)$, after a shift of the integration variable and introducing again $\mu_{\rm CF}$, we write
$I_{\mathrm{F}}\left(\textbf{P},\Omega\right)$ $=$ $\int\!\!\!\frac{d{\mathbf{k}}}{\left(2\pi\right)^{2}}\frac{\Theta\left(-\xi_{{\mathbf{k}}}^{\mathrm{F}}\right)}{{P^{2}\over 2m_{\rm F}}+{({\mathbf{P}}-{\mathbf{k}})^{2}\over 2m_{\rm B}}-\mu_{\rm CF}-i\Omega+{\varepsilon_{0}}}\phantom{aaa}$ (A.12)
$\approx$ $\int\!\!\!\frac{d{\mathbf{k}}}{\left(2\pi\right)^{2}}\frac{\Theta\left(-\xi_{{\mathbf{k}}}^{\mathrm{F}}\right)}{{\varepsilon_{0}}}=\frac{n_{\mu_{\mathrm{F}}}^{0}}{{\varepsilon_{0}}}$ (A.13)
where $n_{\mu_{\mathrm{F}}}^{0}={m_{\mathrm{F}}}{\mu_{\mathrm{F}}}\Theta(\mu_{\rm F})/2\pi$, and, to obtain (A.13), we have neglected all terms in the denominator of Eq. (A.12) except for the large energy scale ${\varepsilon_{0}}$.
@@
Figure 17: Diagrammatic representation of Eq. (B.1) for the 3-body $T$-matrix $T_{3}$. The double line with an arrow or single line with an arrow represents a $T_{2}$ or a vacuum fermionic (F) or bosonic (B) propagator, respectively. Lines without arrows are dangling connection points defining the momentum flow.
We briefly illustrate the evaluation of the 3-body $T$-matrix within the Born approximation in the 2D vacuum. We start by recalling the Skorniakov-Ter-Martirosian equation for the 3-body $T$-matrix $T_{3}(k,p;K)$ (see Fig. 17) describing the scattering between a BF dimer and a fermionic atom of initial three-momenta $K-k$ and $k$ and final three-momenta $K-p$ and $p$, respectively [132, 124]:
$T_{3}(k,p;K)=-G_{\mathrm{B}}^{0}(K-k-p)-\int\!\!\!\frac{d\mathbf{q}}{(2\pi)^{2}}\frac{dq_{0}}{(2\pi)}G_{\mathrm{B}}^{0}(K-k-q)G_{\mathrm{F}}^{0}(q)T_{2}(K-q)T_{3}(q,p;K),$ (B.1)
@@
$G_{s}^{0}(q)=\frac{1}{q_{0}-\mathbf{q}^{2}/2m_{s}+i0^{+}}\quad\quad s=\mathrm{B,F}.$ (B.2)
Within the Born approximation, $T_{3}$ is replaced with the first term on the right-hand side of Eq. (B.1). Moreover, for low energy scattering one can set $k\simeq p\simeq 0$ and $K\equiv\{\textbf{0},-{\varepsilon_{0}}\}$, (taking into account the binding energy ${\varepsilon_{0}}$ of the BF dimer). Recalling that $T_{2}$ is a dimer propagator except for the form factor $2\pi{\varepsilon_{0}}/m_{r}$ (see Eq. (20)) due to its composite nature, the same normalization factor is applied to $T_{3}$, which thus becomes independent of the two-body scattering length $a_{\mathrm{BF}}$
${2\pi{\varepsilon_{0}}\over m_{r}}T_{3}(0,0;\{\textbf{0},-{\varepsilon_{0}}\})\simeq-{2\pi{\varepsilon_{0}}\over m_{r}}G_{\mathrm{B}}^{0}(\textbf{0},-{\varepsilon_{0}})={2\pi\over m_{r}}.$ (B.3)